Solve by completing the square
Find two integers such that the first is 5 more than twice the second and the sum of their squares is 850.

Answer by mananth(16946)

Find two integers such that the first is 5 more than twice the second and the sum of their squares is 850.

Find two integers such that the first is 5 more than twice the second
let second be x
First will be (2x+5)

x^2+(2x+5)^2=850
x^2+4x^2+20x+25=850
5x^2 +20x= 825
/5
x^2+4x=165
x^2+4x+4 = 165+4
(x+2)^2 = 169
square root
x+2 = +/- 13
x= 11 OR x=-15
and the sum of their squares is 850.